OpenStudy (anonymous):

The table below represents a linear function f(x) and the equation represents a function g(x). x f(x) −1 −11 0 −1 1 9 g(x) g(x) = 5x + 1 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points)

OpenStudy (anonymous):

I know the slope is 5

OpenStudy (welshfella):

yes the y intercept is the last number in g(x) = 5x + 1

OpenStudy (welshfella):

slope of g(x) is 5 and y-intercept is +1

OpenStudy (welshfella):

or you could just write it as 1.

OpenStudy (welshfella):

to find the slope of f(x) you need to find the value of difference of y values in 2 consecutive points -------------------------------------- difference in x values in 2 cosecutive points

OpenStudy (welshfella):

first 2 points are ( -1,-11) and (0 -1) so the slope = (1 - (-11)) --------- (0 - (-1))

OpenStudy (welshfella):

remember that 2 negative togethher like in the above become a plus

OpenStudy (welshfella):

example -2-(-4) = - 2 + 4 = 2